OpenStudy (anonymous):

Which of the following is the simplified form of 9 times x cubed over 4 times y to the fourth power divided by the fraction 3 times x to the fifth power over 8 times y squared

OpenStudy (zzr0ck3r):

\[\frac{9x^3}{\frac{4y^4}{\frac{3x^5}{8y^2}}}\]

OpenStudy (zzr0ck3r):

?